Determine the tonicity (iso, hypo, or hypertonic) of the following fluids.

D5W + .02% NaCl

D5W + 40 mEq KCL
Answer #1

The terms tonicity refers to the concentration of IV solution. The IV solutions are divided into three categories according to their tonicity includes, hypertonic, hypotonic, and isotonic. The range of tonicity is measured in milliosmoles, and the normal range is 240 to 340 mOsm: + 50 mOsm and/or − 50 mOsm of 290 mOsm.

1. Hypertonic solutions cause water molecules to diffuse out of the cells and exert a hyper-osmolar effect. Eg: D5 0.9% normal saline(osmolarity : 560 mOsm.).

2.Hypotonic solutions cause water molecules to diffuse into the cells and exert a hypo-osmolar effect. Eg:0.45%NaCl(Osmolarity: 154 mOsm)

3. Isotonic solutions maintain the same concentration of water molecules on both sides of the cell, so no net movement occurs. The osmolarity of isotonic solutions is 240 to 340 mOsm, similar to blood. Eg: Ringer Lactate and 0.9% normal saline solution (NaCl) are the best examples.

Here, discussing the tonicity of two given solutions

D5W + .02% NaCl

The Osmolarity of D5W is 250, which is isotonic solution

The Osmolarity of 0.45%NaCl is 154, therefore

The Osmolarity of 0.02% NaCl (X) = 154/0.45*X/0.02

= 154*0.02/0.45

= 1.386

So, The osmolarity of D5W + .02% NaCl =250+1.386

= 251.386, Which is Isotonic Solution

Even the D5% in 0.2% NaCl is isotonic solution as the osmolarity is 326.

D5W + 40 mEq KCL

The Osmolarity of D5W is 250, which is isotonic solution

The osmolarity of KCL 40mEq /100 ml is 799, However the osmolarity of KCL in which ever solution either in water or in NaCl is going to be hypertonic solution.

Therefore, D5W + 40 mEq KCL will have a osmolarity over the range of 240-340 mOsm.

So, it can be concluded that the solution will be hypertonic solution.